# Sums of squares of integers from residue classes

### v1

@@ 3. Almost square universality of ${\mathcal{A}}_{d,m}$
Since $\text{ord}_{p}(n)=1$, the integer $n$ is not representable as a sum of two squares of integers. On the other hand, by part (1), to write $n=mMt_{0}+2d^{2}$ as a sum of $r$ squares in ${\mathcal{A}}_{d,m}^{2}$, it is necessary that $r\equiv 2\ (\mathrm{mod}\ Mm)$. Thus $r\geq mM+2$. Since there are infinitely many such $n$, we conclude that $\mathrm{ASU}(\mathcal{A}_{d,m})\geq Mm+2$.
Now suppose $m\equiv 1\ (\mathrm{mod}\ 2)$. Consider a positive integer $n=mt_{0}+3d^{2}$ such that $n\equiv 7\ (\mathrm{mod}\ 8)$. Then $n$ is not representable as a sum of three squares of integers. By part (1), in order to write this $n$ as a sum of $r$ squares in ${\mathcal{A}}_{d,m}^{2}$, it is necessary that $r\geq m+3$. Since there are infinitely many such $n$, we conclude that $\mathrm{ASU}(\mathcal{A}_{d,m})\geq m+3$. ∎
As a direct consequence of Proposition 3.1, we obtain the following corollary concerning $\mathrm{ASU}(\mathcal{A}_{d,m})$.
@@
Let $m\in{\mathbb{N}}$ and $d\in{\mathbb{Z}}$ be such that $(m,d)=1$. Then
$\mathrm{ASU}(\mathcal{A}_{d,m})=\begin{cases}m+3&\text{if }m\equiv 1\ (\mathrm{mod}\ 2),\\ Mm+2\text{ or }Mm+3&\text{if }m\equiv 0\ (\mathrm{mod}\ 2).\end{cases}$
Moreover, for $m\equiv 0\ (\mathrm{mod}\ 2)$, we have $\mathrm{ASU}(\mathcal{A}_{d,m})=Mm+2$ if and only if every sufficiently large integer $n\equiv 3d^{2}\ (\mathrm{mod}\ m)$ can be written as a sum of three squares in ${\mathcal{A}}_{d,m}^{2}$.
According to Corollary 3.2, determining $\mathrm{ASU}(\mathcal{A}_{d,m})$ reduces to the case when $m$ is even and to expressing every sufficiently large integer $n\equiv 3d^{2}\ (\mathrm{mod}\ m)$ as a sum of three squares in ${\mathcal{A}}_{d,m}$. We address this problem by studying representations of integers by ternary lattice cosets.
@@
$X_{2}=L_{2}+\nu=\frac{1}{m}(e_{1}+e_{2}+e_{3})=L_{2}+\frac{1}{2}(e_{1}+e_{2}+e_{3}).$
By [14, Proposition 2.3], it follows that $O^{+}(X_{2})=O^{+}(L_{2})$, and hence $\theta\left(O^{+}({X_{2}^{d}})\right)={\mathbb{Z}}_{2}({\mathbb{Q}}_{2}^{\times})^{2}$ as described in the proof of [1, Proposition 4.2].
@@
as in [11, (2.1) and (2.2)]. By setting $m^{\prime}=m+2$, we observe that $K+\mu=L+\left(\frac{m-2}{2}\right)\nu$, and $(m,\frac{m-2}{2})=1$ since $m\equiv 0\ (\mathrm{mod}\ 4)$. Therefore, for all $d$ with $(m,d)=1$, we have $\theta\left(O^{+}({X_{p}^{d}})\right)=\theta\left(O^{+}({K+\mu})\right)$. Note also that
@@
When $m\equiv 2\ (\mathrm{mod}\ 4)$, the above computations give
$\theta\left(O^{+}({X_{p}^{d}})\right)=\begin{cases}{\mathbb{Z}}_{2}({\mathbb{Q}}_{2}^{\times})^{2}&\text{if }p=2,\\ ({\mathbb{Q}}_{3}^{\times})^{2}&\text{if }p=3\text{ and }3\mid m,\\ {\mathbb{Z}}_{p}^{\times}({\mathbb{Q}}_{p}^{\times})^{2}&\text{otherwise}.\end{cases}$
@@
Hence, by Corollary 3.2, it suffices to show that there are infinitely many odd primes $\ell$ satisfying (3.8) and $3\ell^{2}\equiv 3d^{2}\ (\mathrm{mod}\ m)$, equivalently, $\ell^{2}\equiv d^{2}\ (\mathrm{mod}\ 4m_{0})$. Write $m_{0}=2^{a}3^{b}m_{1}$ with $(m_{1},6)=1$. Then by the Chinese Remainder Theorem and Hensel’s Lemma, there exists $\ell$ such that $\ell\equiv d\ (\mathrm{mod}\ m_{1})$ and
@@
Such $\ell$ satisfies $\ell^{2}\equiv d^{2}\ (\mathrm{mod}\ 4m_{0})$. By Dirichlet’s theorem on arithmetic progression. This concludes the proof that $\mathrm{ASU}({\mathcal{A}}_{d,m})=2m+3$ when $m\equiv 0\ (\mathrm{mod}\ 12)$. ∎
@@ 4. Square universality of ${\mathcal{A}}_{d,m}$
$\sum_{i=1}^{r}(mx_{i}+1)^{2}=(m+1)^{2}+(k-1)\cdot(m-1)^{2}+(r-k)=(k+1)^{2}(m-1)^{2}-(m^{2}-6m+1).$
Hence every integer $n$ in the range
$(k+1)^{2}(m-1)^{2}-(m^{2}-6m+1)\leq n\leq(k+1)^{2}(m-1)^{2}+(4m-k-1)$
@@
Although Proposition 3.1 (2) provides an effective bound for almost universality, it remains somewhat too large to guarantee that every small integer is representable as a sum of squares in ${\mathcal{A}}_{d,m}^{2}$. In the following lemma, we derive a sharper bound, from which Lemma 4.2 suffices to complete the verification of $\mathrm{SU}(\mathcal{A}_{d,m})$ for $d\equiv\pm 1\ (\mathrm{mod}\ m)$.
@@
Let $m\geq 6$ be an integer, and define $M=1$ if $m\equiv 1\ (\mathrm{mod}\ 2)$, $M=4$ if $m\equiv 2\ (\mathrm{mod}\ 4)$, and $M=2$ if $m\equiv 0\ (\mathrm{mod}\ 4)$. Then every integer $n\geq 2m(m-1)^{2}$ can be written as a sum of at most $Mm+3$ squares in ${\mathcal{A}}_{d,m}^{2}$.
@@
Write $n=mt+r_{0}$ with unique integers $t,r_{0}\in{\mathbb{Z}}$ such that $0\leq r_{0}<m$. Let $r$ be the unique integer satisfying $m+4\leq r\leq Mm+3$ and $t\equiv\frac{r-r_{0}}{m}\ (\mathrm{mod}\ M)$. Note that the equation
@@
One may verify that, with $e\in\{1,2\}$ such that $e\equiv m\ (\mathrm{mod}\ 2)$,
@@
    $S\subseteq\left\{\sum_{i=5}^{k}\frac{1}{e}(mx_{i}^{2}+2x_{i})\mid x_{i}\in\{0,\pm 1\}\right\}$,
@@
Choosing $s\in S$ satisfying (2), we set
@@
$\begin{cases}x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}=a,\\ x_{1}+x_{2}+x_{3}+x_{4}=1.\end{cases}$
@@
$\mathrm{SU}({\mathcal{A}}_{d,2})=\mathrm{SU}({\mathcal{A}}_{d,4})=\mathrm{SU}({\mathbb{Z}}-2{\mathbb{Z}})=10\quad\text{and}\quad\mathrm{SU}({\mathcal{A}}_{d,3})=\mathrm{SU}({\mathbb{Z}}-3{\mathbb{Z}})=6\$
@@
If $m=5$ or $m=7$, then every integer $n\geq\frac{25}{4}m^{4}+1$ is a sum of at most $m+3$ squares in ${\mathcal{A}}_{d,m}$ by Proposition 3.1. For $m=7$, a direct check shows that every positive integer $n<\frac{25}{4}\cdot 7^{4}+1=15007.25$ is a sum of at most $m^{2}-2m=35$ squares in ${\mathcal{A}}_{d,m}^{2}$, hence $\mathrm{SU}({\mathcal{A}}_{d,7})=m^{2}-2m$. On the other hand, for $m=5$, all positive integers $n\leq\frac{25}{4}\cdot 5^{4}+1=3907.25$ are sums of at most $m^{2}-2m=15$ squares in ${\mathcal{A}}_{d,m}^{2}$, except for $n=31$. Thus, $\mathrm{SU}({\mathcal{A}}_{d,5})=16$.

### v2

@@ 3. Almost square universality of ${\mathcal{A}}_{d,m}$
Since $\text{ord}_{p}(n)=1$, the integer $n$ cannot be expressed as a sum of two squares of integers. On the other hand, by part (1), writing $n=mMt_{0}+2d^{2}$ as a sum of $r$ squares in ${\mathcal{A}}_{d,m}^{2}$ requires that $r\equiv 2\ (\mathrm{mod}\ Mm)$. Hence $r\geq mM+2$. Since there exist infinitely many such $n$, we conclude that $\mathrm{ASU}(\mathcal{A}_{d,m})\geq Mm+2$.
Now suppose that $m\equiv 1\ (\mathrm{mod}\ 2)$. We construct infinitely many positive integers $n_{j}=mt_{j}+4d^{2}$ for $j\geq 0$ that cannot be expressed as a sum of four squares in ${\mathcal{A}}_{d,m}^{2}$. By part (1), expressing $n_{j}$ as a sum of elements of ${\mathcal{A}}_{d,m}^{2}$ requires at least $m+4$ squares, which implies that $\mathrm{ASU}(\mathcal{A}_{d,m})\geq m+4$.
We first choose such an integer $n_{0}$ with $t_{0}\in 2{\mathbb{Z}}$. We begin with the case when $0<d\leq\frac{m-1}{2}$. Note that the three smallest squares in ${\mathcal{A}}_{d,m}^{2}$ are $d^{2}$, $(-m+d)^{2}$, $(m+d)^{2}$, listed in increasing order. Consequently, the smallest two integers and the next several integers that can be expressed as sums of four squares in ${\mathcal{A}}_{d,m}^{2}$ are
$4d^{2}$, $3d^{2}+(-m+d)^{2}$, $2d^{2}+2(-m+d)^{2}$, $d^{2}+3(-m+d)^{2}$, $4(-m+d)^{2}$, and $3d^{2}+(m+d)^{2}$.
Writing each of these integers in the form $mt+4d^{2}$, the corresponding values of $t$ are
0, m2dm-2d, 2(m2d)2(m-2d), 3(m2d)3(m-2d), 4(m2d)4(m-2d), and $m+2d$.
If $d<\frac{m-1}{2}$, then $m-2d\geq 3$, and hence we may take $t_{0}=2$. If $d=\frac{m-1}{2}$, then $m+2d=2m-1\geq 9$, so we may take $t_{0}=6$.
Next, consider the case $\frac{m+1}{2}\leq d\leq m-1$. Note that the three smallest squares in ${\mathcal{A}}_{d,m}^{2}$ are $(-m+d)^{2}$, $d^{2}$, $(-2m+d)^{2}$, listed in increasing order. Hence, the smallest two and the next several integers that can be expressed as sums of four squares in ${\mathcal{A}}_{d,m}^{2}$ are
$4(-m+d)^{2}$, $3(-m+d)^{2}+d^{2}$, $2(-m+d)^{2}+2d^{2}$, $(-m+d)^{2}+3d^{2}$, $4d^{2}$, and $3(-m+d)^{2}+(-2m+d)^{2}$.
Writing each of these integers as $mt+4d^{2}$, the corresponding values of $t$ are
4(m2d)4(m-2d), 3(m2d)3(m-2d), 2(m2d)2(m-2d), m2dm-2d, 0, and 7m10d7m-10d.
If $d>\frac{m+1}{2}$, then $m-2d\leq-3$, and hence we may choose $t_{0}$ to be an even integer strictly between 4(m2d)4(m-2d) and 3(m2d)3(m-2d). If $d=\frac{m+1}{2}$, then $7m-10d=2m-5\geq 5$, and we may take $t_{0}=2$.
For each $j\in{\mathbb{N}}$, we define
$t_{j}:=\frac{4^{\ell j}(mt_{0}+4d^{2})-4d^{2}}{m}\quad\text{and}\quad n_{j}:=mt_{j}+4d^{2}=4^{\ell j}(mt_{0}+4d^{2}),$
where $\ell:=\text{ord}_{m}(2)=\text{ord}_{m}\left(\frac{m+1}{2}\right)$. Since the numerator of $t_{j}$ is a multiple of $m$, we have $t_{j}\in{\mathbb{N}}$.
We claim that $n_{j}$ cannot be written as a sum of four squares in ${\mathcal{A}}_{d,m}^{2}$ for any $j$. Assume to the contrary that
$n_{j}=\sum_{i=1}^{4}(mx_{i}+d)^{2}$
for some $x_{i}\in{\mathbb{Z}}$. Since $n_{j}$ is a multiple of $8$ (as $t_{0}$ is even), it follows that $mx_{i}+d\equiv 0\ (\mathrm{mod}\ 2)$ for all $i$. Noting that $\frac{mx_{i}+d}{2}\equiv d\cdot\frac{m+1}{2}\ (\mathrm{mod}\ m)$, we have
$4^{\ell j-1}(mt_{0}+4d^{2})=\sum_{i=1}^{4}\left(my_{i}+d\cdot\frac{m+1}{2}\right)^{2}$
for some $y_{i}\in{\mathbb{Z}}$. Iterating this argument $\ell-1$ more times, we obtain
$n_{j-1}=4^{\ell(j-1)}(mt_{0}+4d^{2})=\sum_{i=1}^{4}\left(mz_{i}+d\cdot\left(\frac{m+1}{2}\right)^{\ell}\right)^{2}$
for some $z_{i}\in{\mathbb{Z}}$. Since $\text{ord}_{m}\left(\frac{m+1}{2}\right)=\ell$, we have
$n_{j-1}=\sum_{i=1}^{4}\left(mX_{i}+d\right)^{2}$
for some $X_{i}\in{\mathbb{Z}}$. Repeating this procedure, we eventually obtain a representation of $n_{0}$ as a sum of four squares in ${\mathcal{A}}_{d,m}^{2}$, contradicting the choice of $n_{0}$. This proves the claim and completes the proof. ∎
As a direct consequence of Proposition 3.1 and from the fact that $\mathrm{ASU}({\mathcal{A}}_{d,3})=\mathrm{ASU}({\mathbb{Z}}\setminus 3{\mathbb{Z}})=6$ for any $d\in{\mathbb{Z}}$ with $(d,3)=1$ by [15], we obtain the following corollary concerning $\mathrm{ASU}(\mathcal{A}_{d,m})$.
@@
Let $m\in{\mathbb{N}}$ and $d\in{\mathbb{Z}}$ be such that $(m,d)=1$, and define $M=4$ if $m\equiv 2\ (\mathrm{mod}\ 4)$ and $M=2$ if $m\equiv 0\ (\mathrm{mod}\ 4)$. Then
$\mathrm{ASU}(\mathcal{A}_{d,m})=\begin{cases}6&\text{if }m=3,\\ m+4&\text{if }m\equiv 1\ (\mathrm{mod}\ 2),\\ Mm+2\text{ or }Mm+3&\text{if }m\equiv 0\ (\mathrm{mod}\ 2).\end{cases}$
Moreover, for $m\equiv 0\ (\mathrm{mod}\ 2)$, we have $\mathrm{ASU}(\mathcal{A}_{d,m})=Mm+2$ if and only if every sufficiently large integer $n\equiv 3d^{2}\ (\mathrm{mod}\ Mm)$ can be written as a sum of three squares in ${\mathcal{A}}_{d,m}^{2}$.
According to Corollary 3.2, determining $\mathrm{ASU}(\mathcal{A}_{d,m})$ reduces to the case when $m$ is even and to expressing every sufficiently large integer $n\equiv 3d^{2}\ (\mathrm{mod}\ Mm)$ as a sum of three squares in ${\mathcal{A}}_{d,m}^{2}$. We address this problem by studying representations of integers by ternary lattice cosets.
@@
$X_{2}^{d}=L_{2}+\nu=L_{2}+\frac{1}{m}(e_{1}+e_{2}+e_{3})=L_{2}+\frac{1}{2}(e_{1}+e_{2}+e_{3}).$
By [14, Proposition 2.3], it follows that $O^{+}(X_{2}^{d})=O^{+}(L_{2})$, and hence $\theta\left(O^{+}({X_{2}^{d}})\right)={\mathbb{Z}}_{2}({\mathbb{Q}}_{2}^{\times})^{2}$ as described in the proof of [1, Proposition 4.2].
@@
as in [11, (2.1) and (2.2)]. By setting $m^{\prime}=m+2$, we observe that $K+\mu=L+\left(\frac{m-2}{2}\right)\nu$, and $(m,\frac{m-2}{2})=1$ since $m\equiv 0\ (\mathrm{mod}\ 4)$. Therefore, for all $d$ with $(m,d)=1$, we have $\theta\left(O^{+}({X_{p}^{d}})\right)=\theta\left(O^{+}({K_{p}+\mu})\right)$. Note also that
@@
Now we consider the case when $m\equiv 2\ (\mathrm{mod}\ 4)$. Since the computation of $\theta(O^{+}(X^{d}_{p}))$ for odd primes $p$ in [11, Proposition 3.4] applies unchanged in this setting, the above arguments yield
$\theta\left(O^{+}({X_{p}^{d}})\right)=\begin{cases}{\mathbb{Q}}_{2}\setminus\{0\}&\text{if }p=2,\\ ({\mathbb{Q}}_{3}^{\times})^{2}&\text{if }p=3\text{ and }3\mid m,\\ {\mathbb{Z}}_{p}^{\times}({\mathbb{Q}}_{p}^{\times})^{2}&\text{otherwise}.\end{cases}$
@@
Hence, by Corollary 3.2, it suffices to show that there are infinitely many odd primes $\ell$ satisfying (3.10) and $3\ell^{2}\equiv 3d^{2}\ (\mathrm{mod}\ Mm)$, equivalently, $\ell^{2}\equiv d^{2}\ (\mathrm{mod}\ 4Mm_{0})$, where $M=4$ if $m\equiv 2\ (\mathrm{mod}\ 4)$ and $M=2$ if $m\equiv 0\ (\mathrm{mod}\ 4)$. Write $Mm_{0}=2^{a}3^{b}m_{1}$ with $(m_{1},6)=1$. Then by the Chinese Remainder Theorem and Hensel’s Lemma, there exists $\ell$ such that $\ell\equiv d\ (\mathrm{mod}\ m_{1})$ and
@@
Such $\ell$ satisfies $\ell^{2}\equiv d^{2}\ (\mathrm{mod}\ 4Mm_{0})$. By Dirichlet’s theorem on arithmetic progression. This concludes the proof that $\mathrm{ASU}({\mathcal{A}}_{d,m})=2m+3$ when $m\equiv 0\ (\mathrm{mod}\ 12)$. ∎
@@ 4. Square universality of ${\mathcal{A}}_{d,m}$
$\sum_{i=1}^{r}(mx_{i}+1)^{2}=(m+1)^{2}+(k-1)\cdot(m-1)^{2}+(r-k)=(k+1)\cdot(m-1)^{2}-(m^{2}-6m+1)+(r-k).$
Since $k\leq r\leq m^{2}-2m$, every integer $n$ in the range
$(k+1)\cdot(m-1)^{2}-(m^{2}-6m+1)\leq n\leq(k+1)\cdot(m-1)^{2}+(4m-k-1)$
@@
Although Proposition 3.1 (2), (3) provide effective bounds for almost universality, it remains somewhat too large to guarantee that every small integer is representable as a sum of squares in ${\mathcal{A}}_{d,m}^{2}$. In the following lemma, we derive a sharper bound, from which Lemma 4.2 suffices to complete the verification of $\mathrm{SU}(\mathcal{A}_{d,m})$ for $d\equiv\pm 1\ (\mathrm{mod}\ m)$.
@@
Let $m\geq 6$ be an integer, $e\in\{1,2\}$ be such that $e\equiv m\ (\mathrm{mod}\ 2)$, and define $M=1$ if $m\equiv 1\ (\mathrm{mod}\ 2)$, $M=4$ if $m\equiv 2\ (\mathrm{mod}\ 4)$, and $M=2$ if $m\equiv 0\ (\mathrm{mod}\ 4)$. Then every integer $n\geq 2m(m-1)^{2}$ can be written as a sum of at most Mm+5eMm+5-e squares in ${\mathcal{A}}_{d,m}^{2}$.
@@
Write $n=mt+r_{0}$ with unique integers $t,r_{0}\in{\mathbb{Z}}$ such that $0\leq r_{0}<m$. Let $r$ be the unique integer satisfying $m+6-e\leq r\leq Mm+5-e$ and $t\equiv\frac{r-r_{0}}{m}\ (\mathrm{mod}\ M)$. Note that the equation
@@
One may observe that
@@
    $S\subseteq\left\{\sum_{i=5}^{Mm+5-e}\frac{1}{e}(mx_{i}^{2}+2x_{i})\mid x_{i}\in\{0,\pm 1\}\right\}$,
@@
The second property follows since $S$ form a complete system of residues modulo $2\frac{m}{e}$. Choosing $s\in S$ satisfying (2), we set
@@
$x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}=a\quad\text{and}\quad x_{1}+x_{2}+x_{3}+x_{4}=1.$
@@
$\mathrm{SU}({\mathcal{A}}_{d,2})=\mathrm{SU}({\mathcal{A}}_{d,4})=\mathrm{SU}({\mathbb{Z}}\setminus 2{\mathbb{Z}})=10\quad\text{and}\quad\mathrm{SU}({\mathcal{A}}_{d,3})=\mathrm{SU}({\mathbb{Z}}\setminus 3{\mathbb{Z}})=6\$
@@
If $m=5$ or $m=7$, then every integer $n\geq\frac{1}{5}m^{4}+m$ is a sum of at most $m+4$ squares in ${\mathcal{A}}_{d,m}^{2}$ by Proposition 3.1. For $m=7$, a direct check shows that every positive integer $n<\frac{1}{5}\cdot 7^{4}+7=487.2$ is a sum of at most $m^{2}-2m=35$ squares in ${\mathcal{A}}_{d,m}^{2}$, hence $\mathrm{SU}({\mathcal{A}}_{d,7})=m^{2}-2m$. On the other hand, for $m=5$, all positive integers $n\leq\frac{1}{5}\cdot 5^{4}+5=130$ are sums of at most $m^{2}-2m=15$ squares in ${\mathcal{A}}_{d,m}^{2}$, except for $n=31$. Thus, $\mathrm{SU}({\mathcal{A}}_{d,5})=16$.